OpenStudy (anonymous):

Find the amount if $800 is invested at 5% compounded monthly for 10 months.

OpenStudy (ghostgate):

Alright so if you go to a bank put in $800 and it has a 5% interest per month for 10 months here's how you find the answer: (Scroll down to the bottom to get your answer if you don't want to know the entire process to get this answer) First you got to find the number that is 5% of $800. This can be done like so: $800 / 5 = $160 which is 5% of $800. $800 + 16- = $960 for $800 being in the bank for 1 month. Next you got to find the number that is 5% of 960. This can be done like this: $960 / 5 = 192 which is 5% of $960. $960 + 192 = $1,152 for $800 being in the bank for 2 months. Next you got to find the number that is 5% of 1,152 now. This can be done like this: $1,152 / 5 = 230.2 which is 5% of $1,152. $1,152 + 230.4 = $1,382.4 for $800 being in the bank for 3 months. I will speed this up for you now: $1,382.4 / 5 = 276.48(5% of 1,382.4) $800 in for 4 months = $1,658.88 $1,658.88 / 5 = 331.776 $800 in for 5 months = $1,990.656 $1,990.656 / 5 = 398.1312 $800 in for 6 months = $2,388.7872 $2,388.7872 / 5 = 477.75744 $800 in for 7 months = 2,866.54464 $2,866.54464 / 5 = 573.308928 $800 in for 8 months = $3,439.853568 3,439.853568 / 5 = 687.9707136 $800 in for 9 months = $4,127.8242816 $4,127.8242816 / 5 = 825.56485632 Finally if $800 dollars was in for 10 months your answer would be(If my math is correct): $4,953.38913792 Estimated(by tenths): $4,953.4 Estimated(by units): $4,953 dollars. Hope this helps! G'day!

OpenStudy (ghostgate):

Also here's how I broke down the answer. If you want to know how I found 5% of a number so easily take 100 for example and if you divide it by 5 you get 20 and if 20 is multiplied by 5 you get 100. Its a pretty simple way to get your answer. So keep that in mind. Also sorry if this is wrong or for the mistake I made: Instead of putting 160 I put 16- by accident.

OpenStudy (anonymous):

all of these answers are wrong.

OpenStudy (anonymous):

\[F = P(1 + \frac{ r }{ n })^{nt}\] F = Future value after t periods P = Present value (initial investment) r = annual nominal interest rate (not reflecting the compounding) n = number of times the interest is compounded per year t = number of years the money is borrowed for this was copy/pasted from Wikipedia. unless stated otherwise, the interest is assumed yearly. in this case, you would use what mathman said. however this is a monthly rate. \[F = 800(1 + \frac{ 0.05 }{ 12 })^{12*\frac{ 10 }{ 12 }} = $833.97\]

OpenStudy (anonymous):

correcting unclear sentence: unless stated otherwise, the interest rate is assumed to be yearly. in the case where it is yearly, you would use the formula mathman used

OpenStudy (anonymous):

all of these are wrong..

OpenStudy (anonymous):

find he amount after 6 years if $800 is invested at 17% compounded annually

OpenStudy (anonymous):

800*1.17=936*1.17=1095.12*1.17=1281.29*1.17=1499.11*1.17=1753.96*1.17= 2052.13